Hoodmemes:

My mom was at the store with a 15% off coupon. She spent $25.00 with 7% sales tax. How much is the total?

Hoodmemes:

\[7\div100=0.07\] \[25.00+0.07=25.07\] \[15\div100=0.15\] \[25.07-0.15=24.92\] I'm correct right @Convert ?

darkknight:

When it asks for something 15% off or 7% sales tax it asks with regards to the initial amount she spent, which is $25, so you take 7% of 25 not 1 and likewise with 15% of 25 not 1

darkknight:

25.07 is not 7% more than 25

Convert:

@Hoodmemes No, 1st off what you would do is take the 15% off before you add the sales tax too the total which gives you 25(0.15) which is multiplication not subtraction. After you solve this you get $3.75 which you then subtract from the original total: 25-3.75 which gives you $21.25. After you do this you then add the sales tax because you check out after you buy the item: 21.25 (0.07) which gives you 1.4875 which round too $1.49 which you then add to the total: 21.25+1.49= $22.74 which would be the final answer. Answer: $22.74

Hoodmemes:

Oh ok thanks dude

darkknight:

np

Convert:

Happy to help! :)

Convert:

Wait who were you talking too oops lol